OpenStudy (xochital):

What is the simplified form of the expression 4the square root of v. + 4 the square root of v to the sixth power. − the square root of the quantity 25 times v. − the square root of the quantity 9 times v to the sixth power.? 6v2the square root of v. v3 −the square root of v. 6v5the square root of v. the square root of v end root minus v cubed.

OpenStudy (anonymous):

\[4\sqrt{v}+4\sqrt{v^6}-\sqrt{25v}-\sqrt{9 v^6}\]

OpenStudy (anonymous):

is this the ques?

OpenStudy (anonymous):

@xochital

OpenStudy (xochital):

yes :)

OpenStudy (anonymous):